OpenStudy (anonymous):

The distance between the two parallel lines L1: −7 x − 10 y=−4, L2: −7 x − 10 y=3 is ?????

OpenStudy (anonymous):

-7x-10y=-4 can be re-written as: -10y=7x-4. -7x-10y=3 can also be re-written as -10y = 7x +3 i.e. in the form y=mx+c this shows that the slopes are indeed the same (they lines are parallel), and the distance is 7. i.e. from the first c (on the y-axis) to the second c. (i.e. from -4 to 3)

OpenStudy (anonymous):

i tried 7 it didn't wok

OpenStudy (anonymous):

change signs first, they become: 10y=-7x-3 and 10y=-7x + 4 now divide across by 10. y=-7x-3/10 and y= -7x +4/10 so answer is 4/10+ 3/10 = 7/10 or 0.7

OpenStudy (anonymous):

apologies for oversight!

OpenStudy (anonymous):

again says incorrect

OpenStudy (anonymous):

It might say incorrect, but I am certain this is the answer for the equations you gave, so either: a) did you try both values 7/10 AND 0.7 b) can you check to see if the equations themselves are correct as per question if they are then there is either a problem with the question or the answer it has.

OpenStudy (anonymous):

A) yes i tried both B) yes they r correct at leas i copied right :D

OpenStudy (anonymous):

The distance between the two lines is not the gap between the y-intercepts. The distance between the two lines is the distance along a perpendicular line. The slope of the perpendicular is the inverse of the slope of either parallel line.